OpenStudy (anonymous):

A stone is thrown upward from ground level. With what minimum speed should the stone be thrown so as to reach a height of 100 feet?

OpenStudy (anonymous):

2) A falling stone is at a certain instant 204 feet above the ground. 2 seconds later it is only 12 feet above the ground. From what height was it dropped?

OpenStudy (anonymous):

which one are we doing? i think you are supposed to use \[h(t)=-16t^2+v_0t+h_0\]

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

if we are doing the first one, since you are starting at "ground level" \[h_0=0\] so you need to solve \[-16t^2+v_0t\geq 100\] for v0

OpenStudy (anonymous):

using the kinematic formula for motion in 2 dimension: \[v_f^2 = v_i^2 + 2a(y_f-yi)\] Where the height it must reach is 100 feet, and you know g=9.8 m/s^2, the right hand side of equation becomes: \[v_f^2 = v_i^2 + 2(-9.8)(100)\] \[(v_f^2 - v_i^2)= 2(-9.8)(100)\] As the stone reaches max hieght, the velocity is zero \[((0)^2 - v_i^2)= 2(-9.8)(100)\] \[v_f= \sqrt{2(9.8)(100)}\] \[v_i=14\sqrt{10} ft/s\]

OpenStudy (anonymous):

we are doing both aslo

OpenStudy (anonymous):

in other words make sure the equation \[-16t^2+v_0t-100=0\] has a solution. use the quadratic formula and make sure the discriminant is positive

OpenStudy (anonymous):

the answer choices for teh first one are 80, 112, 88, 96 or 78

OpenStudy (anonymous):

you wold get \[t=\frac{-v_0\pm\sqrt{v_0^2+6400}}{-32}\] so you have to make sure that \[v_0^2+6400\geq 0\]

OpenStudy (anonymous):

and therefore \[v_0\geq80\]

OpenStudy (anonymous):

so 80 wud b the answer rite

OpenStudy (anonymous):

what about number 2

OpenStudy (anonymous):

i made a mistake it should be \[v_0^2-6400\geq 0\]

OpenStudy (anonymous):

ok in the second one the stone was dropped, so you know \[v_0=0\] but you do not know \[h_0\]

OpenStudy (anonymous):

so the equation will look like \[h(t)=-16t^2+h_0\]

OpenStudy (anonymous):

you know that for a certain t, \[h(t)=-16t^2+h_0=240\] and that for t + 2 it is \[h(t+2)=-16(t+2)^2+h_0=12\] you have to expand the second one and then you will have two equations involving h0 and t, you can solve for both

OpenStudy (anonymous):

the answer chocies for teh second q are 271, 270, 266, 268, or 269

OpenStudy (anonymous):

im confused...i expanded it adn got -16t^2 +66h +ht +64 +h0....but we dont evn have teh time